OpenStudy (anonymous):

Someone please explain to me how to do this: Find the exact value of sec(11 π / 3) using reference angles. Show all of your work. Do NOT use a calculator.

OpenStudy (jdoe0001):

\(\bf \cfrac{11\pi}{3}\implies \cfrac{6\pi}{3}+\cfrac{5\pi}{3}\implies 2\pi+\cfrac{5\pi}{3}\) what do you think?

OpenStudy (jdoe0001):

can you see the reference angle?